If A=3x2-2x-7 and B=x2+9, then what is their sum, A+B?
Answer by palanisamy(496)

Given, A = 3x^2-2x-7 and B = x^2+9
A+B =3x^2-2x-7+x^2+9
A+B = 4x^2-2x+2